Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

11.00
Let probability of hitting the target = p ⇒ p=1/2
Let n be the minimum number of bombs
According to given condition
1 - (?C?P?(1-P)? + ?C?P¹(1-P)?¹) ≥ 99/100
⇒ 2? ≥ (n+1)100
n=10 ⇒ 2¹? ≥ 1100 Reject
n=11 ⇒ 2¹¹ ≥ 1200 Select

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

Given, mean = np = . and variance = npq = α3⇒q=13   and   p=23

P (X=1)=np1qn−1=4243⇒n (23)1 (13)n−1=4243⇒n=6

P (X=4  or  5)=6C4 (23)4 (13)2+6C5 (23)5 (13)1=1627

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

Mean = 4 = μ = np

Variance = σ2=np (1−p)=43⇒4 (1−p)=43⇒p=23⇒n=6

=6C0 (13)6+6C1 (23)1 (13)6+6C2 (23)2 (13)4=14627

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

 P (A/B)=17⇒P (A∩B)P (B)=17

⇒P (B)=79

P (B/A)=25⇒P (A∩B)P (A)=25

⇒P (A)=518

Now,  P (A'∪B)=1−P (A∪B)+P (B)

Both (S1 ) and (S2 ) are true

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

P (En) = n/36 for n = 1, 2, 3, …., 8

P (A)=Any  possible  sum  of   (1, 2, 3, ........., 8) (=α say)36


α36≥45

∴a≥29

If one of the number from {1, 2, ….8} is left then total ≥ 29 by 3 ways

Similarly by leaving terms more 2 or 3 we get 16 more combinations

∴ Total number of different set a possible is 16 + 3

= 19

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

The required probability

=Area of  Region  PQCAPArea  of  Region  ABCA

=12×8×6−12×2×412×8×6

=56

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a True or False Type Questions as classified in NCERT Exemplar

Since,  the  two  given  events  are  not  related  to  the  same  Sample  space.∴The  sum  of  probabilities  of  two  students  getting  distinction  in  their  final      may  be  1.2Hence,  the  given  statement  is  'True'

Thumbs Up IconUpvote Thumbs Down Icon